If you had 20 rabbits. After six months the number of rabbits had increased to 100. If the number of rabbits increased exponentially, then
How many rabbits will you have at the end of one year? (Hint: consider the general exponential model y=ab^(x))

Answer by rothauserc(4718)

a = 20 and b = 100/20 = 5
:
y = 20(5)^x
:
Note 12/6 = 2
:
y = 20(5)^2 = 500 rabbits
: